### v5

@@ Abstract
Multivalued treatments are commonplace in applications. We explore the use of discrete-valued instruments to control for selection bias in this setting. Our discussion revolves around the concept of targeting: which instruments target which treatments. It allows us to establish conditions under which counterfactual averages and treatment effects are point- or partially-identified for composite complier groups. We illustrate the usefulness of our framework by applying it to data from the Head Start Impact Study. Under a plausible positive selection assumption, we derive informative bounds that suggest less beneficial effects of Head Start expansions than the parametric estimates of Kline and Walters (2016).
@@ Introduction
In the binary/binary model, the analyst can often take for granted that switching on the binary instrument makes treatment (weakly) more likely for all or no observations. This is satisfied under the local average treatment effect (LATE)-monotonicity assumption (e.g., Imbens and Angrist, 1994; Vytlacil, 2002; Heckman and Vytlacil, 2007a). With multiple instrument values and multiple treatments, there may be no natural ordering of instrument or treatment values that would give meaning to the word “monotonicity”. Since Heckman and Pinto (2018) defined an “unordered monotonicity” property, various papers have proposed other definitions of (qualified) monotonicity [^5] See Navjeevan and Pinto (2022) for a detailed analysis of some of these proposals..
@@
Even when some sort of monotonicity holds, there exist several groups of compliers—individuals whose treatment assignment changes with the value of the instrument. The multiplicity of treatments and instruments may give rise to a bewildering number of cases, as existing literature demonstrates. Angrist and Imbens (1995) analyzed two-stage least squares (TSLS) estimation when the treatment takes a finite number of ordered values. Closer to us, Heckman, Urzua, and Vytlacil (2006); Heckman and Vytlacil (2007b); Heckman, Urzua, and Vytlacil (2008) discussed the identification of treatment effects in the presence of discrete-valued instruments when assignment to treatment can be modeled as a discrete choice model. Several recent papers have studied the case of binary treatments with multiple instruments. Mogstad, Torgovitsky, and Walters (2021) and Goff (2024a) analyzed the identifying power of different monotonicity assumptions in this context [^6] Mogstad, Torgovitsky, and Walters (2020) further apply their framework of monotonicity with multiple instruments to marginal treatment effects (e.g., Heckman and Vytlacil, 2001, 2005; Carneiro, Heckman, and Vytlacil, 2011).. Others have studied models with binary instruments and multivalued or continuous treatments. Torgovitsky (2015), D’Haultfoeuille and Février (2015), Huang, Khalil, and Yildiz (2019), Caetano and Escanciano (2021), and Feng (2024) developed identification results for different models.
@@
In a wide-ranging contribution, Heckman and Pinto (2018) derived results on partial identification in discrete-instrument, discrete-treatment models; they also showed how additional identifying assumptions, such as unordered monotonicity, can be applied to shrink the identified set of treatment effects for various complier groups. While their results are very general, they are not as transparent as one would like. Our approach to this issue is different: we seek a parsimonious framework within which we can make constructive progress, and that can still be useful in many applications. In order to reduce the complexity of the problem, we start by imposing an additive random-utility model (ARUM) structure. Under ARUM, the selection into treatment depends on mean values and additive, observation-specific shocks. Some, but not all, ARUM models satisfy the unordered monotonicity property of Heckman and Pinto (2018), which was applied by Pinto (2021) to the Moving to Opportunity program. In many applications, some observations are not treated; in others, another treatment value is particularly salient. We call it the “control”. Under ARUM, each treatment $t$ generates a change in the mean value, relative to the control, that depends on the value $z$ of the instrument. It is natural to speak of an instrument value $z$ targeting a treatment value $t$ when it maximizes this change in mean value. Most of our paper relies on the assumption of strict targeting, which obtains when each instrument only changes the mean values of the treatments it targets. Strict targeting holds for instance in models of imperfect compliance when the cost of non-compliance does not depend on its nature. Some of our results also require one-to-one targeting, where each non-zero instrument targets one treatment only, and each treatment (apart from the control) is targeted by one instrument only.
@@
To illustrate, consider a typical randomized experiment with imperfect compliance: (i) individuals are randomly assigned to a treatment branch $t$ (including a non-treatment option 00 0) based on the instrument value $z$ that they draw; (ii) some individuals self-select into a treatment branch $t^{\prime}$ that they prefer, even though they did not draw the corresponding instrument value $z^{\prime}$. In our terminology, $z$ targets $t$ and $z^{\prime}$ targets $t^{\prime}$. Often this mapping is one-to-one; this is what our one-to-one targeting assumption states. Strict targeting of a treatment branch $t$ is more restrictive, as its name indicates. One way to interpret it in a non-compliance context is that it is equally difficult for an average individual in the population to select into treatment $t$ when she was not assigned to that branch, no matter what the experimenter’s intended treatment branch was. To cite two examples, consider the interventions reported in Angrist, Lang, and Oreopoulos (2009) and in Attanasio, Fernández, Fitzsimons, Grantham-McGregor, Meghir, and Rubio-Codina (2014; 2020). These are 4-way factorial randomized experiments: each subject is randomly assigned to a control group, to receive treatment 1, to receive treatment 2, or to receive both treatments. By definition, this is one-to-one targeting. Compliance was very imperfect in Angrist, Lang, and Oreopoulos (2009), and it is described as “high” in the other two papers. If subjects self-selected into treatments on the basis of their expected benefits, then strict targeting is a natural assumption.
@@
Combining ARUM and assumptions on targeting allows us to point-identify the size of some complier groups and the corresponding counterfactual averages and treatment effects on any function of the outcomes, and to partially identify others. We use two examples to demonstrate the identification power and implications of ARUM and targeting. Our first example is a $2\times T$ model where a binary instrument targets only one of $T\geq 3$ treatment values, as in Kline and Walters (2016). In our second example, three unordered treatment values target three instrument values. This $3\times 3$ model was also studied by Kirkeboen, Leuven, and Mogstad (2016) [^9] See also more recent work by Bhuller and Sigstad (2024), Heinesen, Hvid, Kirkeboen, Leuven, and Mogstad (2022), and Nibbering, Oosterveen, and Silva (2022).. Unlike them, we do not assume that the data contains information on next-best alternatives. Whereas the $2\times T$ model satisfies unordered monotonicity under our strongest targeting assumptions, the $3\times 3$ model does not101010 It does satisfy the weaker generalized monotonicity assumption of Bai, Huang, Moon, Shaikh, and Vytlacil (2024), however..
@@
We obtain novel identification results for both examples; they lead to new estimands or bounds for average treatment effects on various groups. Additional identifying assumptions can refine these bounds. One example is what we call positive selection. This assumes that the average outcome for a given treatment $t$ is larger for some response group than for another. Consider for instance the binary instrument case. It seems natural to assume that the always-takers of a treatment get more from it than compliers who only take it if they are incentivized to do so. Positive selection also obtains under weak assumptions in the generalized Roy model. More generally, let us return to our earlier illustration of a randomized experiment under imperfect compliance. Consider the response group of individuals who would end up in treatment $t^{\prime}$ both when drawing $z$ and when drawing $z^{\prime}$. We would expect this response group to have better outcomes under $t^{\prime}$, on average, than the response group that exhibits perfect compliance to $z$ and $z^{\prime}$ draws—assuming that these two response groups end up in the same treatment branches for all other instrument values. This falls exactly under our positive selection assumption. It adds identifying power in both of our leading examples.
@@
To illustrate the usefulness of our framework, we apply it to the Head Start Impact Study (HSIS), a randomized experiment that sought to evaluate the value added of Head Start preschools. Kline and Walters (2016) revisited the HSIS; they took into account the presence of a substitute treatment (alternative preschools in this case). They found that Head Start was only beneficial for children who would not have attended an other preschool program instead. We confirm the importance of taking into consideration alternative preschools when evaluating Head Start. Unlike Kline and Walters (2016), we do not rely on parametric selection models. Under a plausible positive selection assumption, our estimates suggest that the large difference between complier groups that they find can only be rationalized under negative selection into Head Start. As a by-product, we provide an upper bound on the welfare effect of expanding access to Head Start. Interestingly, the estimated upper bound turns out to be lower than the point estimate of Kline and Walters (2016); and it yields a lower marginal value for public funds used in expanding access to Head Start.
@@
The paper is organized as follows. Section 1 defines our framework. In Section 2, we define and discuss the concepts of targeting, one-to-one targeting, and strict targeting. Section 3 derives their implications for the identification of population shares, counterfactual averages, and the effects of the treatments on various complier groups; it also defines and illustrates positive selection. Finally, we present estimation results for Head Start in Section 4. The Appendices contain the proofs of all propositions and lemmata, along with some additional material.
@@ 1 The Framework
We focus in this paper on treatment variables that take discrete values, which we label $t\in\mathcal{T}$. For simplicity, we will call $T=t$ “treatment $t$”. These values do not have to be ordered; e.g., when $t=2$ is available, it does not necessarily indicate “more treatment” than $t=1$. We assume that the only available instruments are discrete-valued, and we label their values as $z\in\mathcal{Z}$.
We will use the standard counterfactual notation: $T_{i}(z)$ and $Y_{i}(t,z)$ denote respectively potential treatments and outcomes. ${\rm 1\kern-3.99994pt1}(A)$ denotes the indicator of set $A$.
The validity of the instruments requires the usual exclusion and independence restrictions:
Assumption 1 (Valid Instruments).
  • (i)
    $Y_{i}(t,z)=Y_{i}(t)$ for all $(t,z)$ in $\mathcal{T}\times\mathcal{Z}$.
  • (ii)
    $Y_{i}(t)$ and $T_{i}(z)$ are independent of $Z_{i}$ for all $(t,z)$ in $\mathcal{T}\times\mathcal{Z}$.
Under Assumption 1, we define $T_{i}:=T_{i}(Z_{i})$ and $Y_{i}:=Y_{i}(T_{i})$.
Throughout the paper, we assume that we observe $(Y_{i},T_{i},Z_{i})$ for each $i$.
@@ 1.1 Restricting Heterogeneity
As in most of this literature, we will need an assumption that restricts the heterogeneity in the counterfactual mappings $T_{i}(z)$. In the binary/binary model, this is most often done by imposing LATE-monotonicity. As is well-known, LATE-monotonicity imposes that (denoting instrument values as $z=0,1$) (i) or (ii) must hold:
@@
for some real numbers $U_{z}(t)$ which are common across observations, and random vectors ${(u_{it})}_{t\in\mathcal{T}}$ that are distributed independently of $Z_{i}$. We do not restrict the codependence of the random variables $u_{it}$. The usual models of multinomial choice belong to this family. ARUM also includes ordered treatments, for which $u_{it}\equiv\sigma(t)u_{i}$ for some increasing positive function $\sigma$.
@@
This inequality in turn is incompatible with the existence of an observation $j$ that has treatment values $t^{\prime}$ under $z$ and $t$ under $z^{\prime}$. Thus we rule out “direct two-way flows”: if a change in the value of an instrument causes an observation to shift from a treatment value $t$ to a treatment value $t^{\prime}$, it can cause no other observation to switch from $t^{\prime}$ to $t$. The argument above is a special case of the general discussion in Heckman and Pinto (2018); their Theorem T-3 shows that the treatment assignment models that satisfy unordered monotonicity for each pair of instrument values can be represented as an ARUM. Not all ARUM models satisfy unordered monotonicity, however; unordered monotonicity excludes a more general class of two-way flows. We will illustrate this point on one of our leading examples in Section 3.3.
@@ 1.2 Assignment to Treatment
    where the vector $(u_{it})_{t\in\mathcal{T}}$ is distributed independently of $Z_{i}$ and has an absolutely continuous distribution with full support on $\mathbb{R}^{|\mathcal{T}|}$.
@@
If we add a third treatment value so that $\mathcal{T}=\{0,1,2\}$, the ARUM assumption starts to bite as it excludes direct two-way flows in the treatment model. However, the combination of Assumptions 1 and 2 is far from sufficient to identify interesting treatment effects in general. In order to better understand what is needed, we now resort to the notion of response-groups of observations, whose members share the same mapping from instruments $z$ to treatments $t$. We first state a general definition [^11] This is analogous to the definitions in Heckman and Pinto (2018)..
@@
To illustrate, consider the binary instrument/binary treatment case. It has a priori $4\displaystyle 2^{2}=$ 4 response vectors, $R\in\{00,01,10,11\}$ with corresponding response-groups $C_{00},C_{01},C_{10},C_{11}$. In this notation, the first number refers to a treatment value with $z=0$ and the second number with $z=1$. For instance, $C_{01}$ refers to those with $T_{i}(0)=0$ and $T_{i}(1)=1$, while the composite response-group $C_{\ast 1}$, for which $R(0)=\{0,1\}$, represents the union of $C_{01}$ and $C_{11}$ The LATE-monotonicity assumption implies that either $C_{01}$ or $C_{10}$ is empty.
@@ 2 Targeting
We start by introducing additional assumptions on the underlying treatment model. We will illustrate these assumptions on three examples: the “binary instrument model” or the “$2\times T$” model; the “$3\times 3$ model”; and a generalized Roy model. We first define them briefly.
@@
Suppose that agents choose occupations $t=0,\ldots,\lvert\mathcal{T}\rvert-1$ on the basis of their expected wages $w_{i}(t)=\bar{w}(t)+\eta_{it}$, net of labor disutilities that depend on the values of the instruments:
$T_{i}(z)=\arg\max_{t=0,\ldots,\lvert\mathcal{T}\rvert-1}(w_{i}(t)-d_{i}(z,t))$
where $d_{i}(z,t)=\bar{d}_{z}(t)+v_{it}$. Potential wages are $Y_{i}(t)=w_{i}(t)+\varepsilon_{it}$. We observe $Z_{i}$, the chosen occupation $T_{i}=T_{i}(Z_{i})$, and realized wages $Y_{i}=Y_{i}(T_{i})$. If the vector of variables $\{(\eta_{it},v_{it})\}_{t=0}^{\lvert\mathcal{T}\rvert-1}$ is independent of $Z_{i}$, this is an ARUM model with $U_{z}(t)=\bar{w}(t)-\bar{d}_{z}(t)$ and $u_{it}=\eta_{it}-v_{it}$.
[…]
@@
Under Assumption 2, assignment to treatment is governed by the differences in mean values $(U_{z}(t)-U_{z}(\tau))$ and by the differences in unobservables $u_{it}-u_{i\tau}$. Only the former depend on the instrument. From now on, we assume that there is a reference treatment value $t_{0}$ whose mean utility does not depend does not depend on the value of the instrument:
@@
In many applications, $t=0$ is a “no-treatment” value, and instruments only change the mean utilities of the other treatments. For instance, tuition subsidies, investment credits, and invitations to training programs have no effect for those who do not attend college, do not invest, or choose not to train. Assumption 3 seems natural in such cases [^12] In the generalized Roy model (Example 3), it holds if the disutility of occupation 0 does not depend on the values of the instruments.. For a counter-example, consider a program of unconditional cash transfers with different values $z$, for which we observe the purchases $t$ of several categories of goods the following month. If a household decides to save the transfer ($t=0$), its mean (discounted) utility will still depend on the value of the transfer $z$ that it received [^13] In that case one could define targeting with the function $\tilde{U}_{z}(t)=U_{z}(t)-U_{z}(0)$. We have not explored the consequences of this alternative definition..
@@
denote the maximum value of $U_{z}(t)$ over $z\in\mathcal{Z}$ and the set of maximizers, respectively. If $Z^{\ast}(t)$ is not all of $\mathcal{Z}$, then we will say that the instrument values $z\in Z^{\ast}(t)$ target treatment value $t$; and we write $t\in T^{\ast}(z)$. We denote by $\mathcal{T}^{\ast}$ the set of targeted treatments and $\mathcal{Z}^{\ast}=\operatorname*{{\bigcup}}_{t\in\mathcal{T}^{\ast}}\;Z^{\ast}(t)$ the set of targeting instruments.

### v6

@@ Abstract
Multivalued treatments are commonplace in applications. We explore the use of discrete-valued instruments to control for selection bias in this setting. Our discussion revolves around the concept of targeting: which instruments target which treatments. It allows us to establish conditions under which counterfactual averages and treatment effects are point- or partially-identified for composite complier groups. We explore the additional identifying power of a positive selection assumption. We illustrate its usefulness by revisiting the findings of Kline and Walters (2016) on the Head Start Impact Study. We derive informative bounds that suggest less beneficial effects of Head Start expansions than their parametric estimates.
@@ Introduction
In the binary/binary model, the analyst can often take for granted that switching on the binary instrument makes treatment (weakly) more likely for all or no observations. This is satisfied under the local average treatment effect (LATE)-monotonicity assumption (e.g., Imbens and Angrist, 1994; Vytlacil, 2002; Heckman and Vytlacil, 2007a). With multiple instrument values and multiple treatments, there may be no natural ordering of instrument or treatment values that would give meaning to the word “monotonicity”. Heckman and Pinto (2018) defined an “unordered monotonicity” property; various papers have proposed other definitions of (qualified) monotonicity [^5] See Navjeevan and Pinto (2022) for a detailed analysis of some of these proposals.. Even when such a condition holds, there exist several groups of compliers—individuals whose treatment assignment changes with the value of the instrument. Since this may give rise to a multiplicity of cases, existing literature has often added assumptions that reduce this complexity. Angrist and Imbens (1995) analyzed two-stage least squares (TSLS) estimation when the treatment takes a finite number of ordered values. Several recent papers have studied the case of binary treatments with multiple instruments. Mogstad, Torgovitsky, and Walters (2021) and Goff (2024a) analyzed the identifying power of different monotonicity assumptions in this context [^6] Mogstad, Torgovitsky, and Walters (2024) further apply their framework of monotonicity with multiple instruments to marginal treatment effects (e.g., Heckman and Vytlacil, 2001, 2005; Carneiro, Heckman, and Vytlacil, 2011).. Others have studied models with binary instruments and multivalued or continuous treatments. Torgovitsky (2015), D’Haultfoeuille and Février (2015), Huang, Khalil, and Yildiz (2019), Caetano and Escanciano (2021), Feng (2024), and Angrist, Santos, and Tecchio (2025) developed identification results for different models.
@@
In a wide-ranging contribution, Heckman and Pinto (2018) derived results on partial identification in discrete-instrument, discrete-treatment models; they also showed how additional identifying assumptions, such as unordered monotonicity, can be applied to shrink the identified set of treatment effects for various complier groups. While their results are very general, they are not as transparent as one would like. Our approach to this issue is different: we seek a parsimonious framework within which we can make constructive progress, and that can still be useful in many applications. In order to reduce the complexity of the problem, we start by imposing an additive random-utility model (ARUM) structure, as did Heckman, Urzua, and Vytlacil (2006, 2008) and Heckman and Vytlacil (2007b). Under ARUM, the selection into treatment depends on mean values and additive, observation-specific shocks. Some, but not all, ARUM models satisfy the unordered monotonicity property of Heckman and Pinto (2018), which was applied by Pinto (2022) to the Moving to Opportunity program.
In many applications, a value of the treatment is especially salient; since it often is the no-treatment value $t=0$, we call it the “control”. Under ARUM, each treatment $t$ generates a change in the mean value, relative to the control, that depends on the value $z$ of the instrument. It is natural to speak of an instrument value $z$ targeting a treatment value $t$ when it maximizes this change in mean value. Most of our paper relies on the assumption of strict targeting, which obtains when each instrument only changes the mean values of the treatments it targets. Strict targeting holds for instance in models of imperfect compliance when the cost of non-compliance does not depend on its nature. Some of our results also require one-to-one targeting, where each non-zero instrument targets one treatment only, and each treatment (apart from the control) is targeted by one instrument only. Finally, we speak of universal targeting when both strict and one-to-one targeting apply.
As we will see, each of these targeting assumptions generate testable implications. These tests can also be used to match instrument values and the treatment values that they target. They only rely on estimates of the generalized propensity scores, which are directly identifiable from the data. In particular, the testable implications of universal targeting that we derive are identical to those of Bai, Huang, and Tabord-Meehan (2025) and are therefore sharp.
@@
To illustrate the usefulness of our framework, we apply it to the Head Start Impact Study (HSIS), a randomized experiment that sought to evaluate the value added of Head Start preschools. Kline and Walters (2016) revisited the HSIS; they took into account the presence of a substitute treatment (alternative preschools in this case). They found that Head Start was only beneficial for children who would not have attended another preschool program instead. In this study, the instrument is binary: a child is offered admission to Head Start or not. Treatment is ternary, as the child may end up in Head Start, in an alternative preschool, or not be enrolled in preschool. In our language, “no preschool” is the reference treatment. Head Start offers of course target Head Start; and since the instrument is binary, targeting is trivially strict. This makes it an example of universal targeting.
@@
As another example, consider a randomized experiment with imperfect compliance. Individuals are assigned to a treatment arm via instrument (targeting), but may self-select into an alternative treatment (imperfect compliance). When random assignments map uniquely to treatments, the design satisfies our one-to-one targeting assumption. Strict targeting further implies that selecting into treatment is equally costly for all individuals not assigned to it. Consider for instance the interventions reported in Angrist, Lang, and Oreopoulos (2009) and in Attanasio, Fernández, Fitzsimons, Grantham-McGregor, Meghir, and Rubio-Codina (2014; 2020). These are 4-way factorial randomized experiments: each subject is randomly assigned to a control group, to receive treatment 1, to receive treatment 2, or to receive both treatments. By definition, this is one-to-one targeting. Compliance was very imperfect in Angrist, Lang, and Oreopoulos (2009), and it is described as “high” in the other two papers. If subjects self-selected into treatments on the basis of their expected benefits, then strict targeting is a natural assumption.
@@
Combining ARUM and assumptions on targeting allows us to point-identify the size of some complier groups and the corresponding counterfactual averages and treatment effects on the outcomes, and to partially identify others. We use two examples to demonstrate the identification power and implications of ARUM and targeting. Our first example is a $2\times T$ model where a binary instrument targets only one of $T\geq 3$ treatment values, as in Kline and Walters (2016). In our second example, three unordered treatment values target three instrument values. This $3\times 3$ model was also studied by Kirkeboen, Leuven, and Mogstad (2016) [^9] See also more recent work by Bhuller and Sigstad (2024), Heinesen, Hvid, Kirkeboen, Leuven, and Mogstad (2025), and Nibbering, Oosterveen, and Silva (2022).. Unlike them, we do not assume that the data contains information on next-best alternatives. Whereas the $2\times T$ model satisfies unordered monotonicity under our strongest targeting assumptions, the $3\times 3$ model does not [^10] It does satisfy the weaker generalized monotonicity assumption of Bai, Huang, Moon, Shaikh, and Vytlacil (2024), however..
@@
We obtain novel identification results for both examples; they lead to new estimands or bounds for average treatment effects on various groups. Additional identifying assumptions can refine these bounds. Our leading example is what we call positive selection. This assumes that the average outcome for a given treatment $t$ is larger for some response group than for another. Consider for instance the binary instrument case. It seems natural to assume that the always-takers of a treatment get more from it than compliers who only take it if they are incentivized to do so. Positive selection also obtains under weak assumptions in the generalized Roy model. More generally, let us return to our earlier illustration of a randomized experiment under imperfect compliance. Consider the response group of individuals who would end up in treatment $t^{\prime}$ both when drawing $z$ and when drawing $z^{\prime}$. We would expect this response group to have better outcomes under $t^{\prime}$, on average, than the response group that exhibits perfect compliance to $z$ and $z^{\prime}$ draws—assuming that these two response groups end up in the same treatment branches for all other instrument values. This falls exactly under our positive selection assumption. It adds identifying power in both of our leading examples.
@@
To show the value of our approach, we apply it to the reanalysis of Head Start by Kline and Walters (2016). We confirm the importance of taking into consideration alternative preschools when evaluating Head Start. Unlike Kline and Walters (2016), we do not rely on parametric selection models. Under a plausible positive selection assumption, our estimates suggest that the large difference between complier groups that they find can only be rationalized under negative selection into Head Start. As a by-product, we provide an upper bound on the welfare effect of expanding access to Head Start. Interestingly, the estimated upper bound turns out to be lower than the point estimate of Kline and Walters (2016); and it yields a lower marginal value for public funds used in expanding access to Head Start.
@@
The paper is organized as follows. Section˜1 defines our framework. In Section˜2, we define and discuss the concepts of targeting, one-to-one targeting, and strict targeting. Section 3 derives their implications for the identification of population shares, counterfactual averages, and the effects of the treatments on various complier groups; it also defines and illustrates positive selection. Finally, we present estimation results for Head Start in Section˜4. The Appendices contain the proofs of all propositions and lemmata, along with some additional material.
@@ 1 The Framework
We focus in this paper on treatment variables that take discrete values, which we label $t\in\mathcal{T}$. For simplicity, we will call $T=t$ “treatment $t$”. These values do not have to be ordered; e.g., when $t=2$ is available, it does not necessarily indicate “more treatment” than $t=1$. We assume that the only available instruments are discrete-valued, and we label their values as $z\in\mathcal{Z}$. Finally, for any $z\in\mathcal{Z}$ and $t\in\mathcal{T}$, we denote the generalized propensity score by
$P(t|z)\equiv\Pr(T_{i}=t|Z_{i}=z).$
@@ 1.1 Restricting Heterogeneity
As in most of this literature, we will need an assumption that restricts the heterogeneity in the counterfactual mappings $T_{i}(z)$, that is, potential treatments. In the binary/binary model, this is most often done by imposing LATE-monotonicity. As is well-known, LATE-monotonicity imposes that (denoting instrument values as $z=0,1$) (i) or (ii) must hold:
@@
for some real numbers $U_{z}(t)$ which are common across observations, and random vectors ${(u_{it})}_{t\in\mathcal{T}}$ that are distributed independently of $Z_{i}$. We do not restrict the codependence of the random variables $u_{it}$.
@@
This inequality in turn is incompatible with the existence of an observation $j$ that has treatment values $t^{\prime}$ under $z$ and $t$ under $z^{\prime}$. Thus we rule out “direct two-way flows”: if a change in the value of an instrument causes an observation to shift from a treatment value $t$ to a treatment value $t^{\prime}$, it can cause no other observation to switch from $t^{\prime}$ to $t$. The argument above is a special case of the general discussion in Heckman and Pinto (2018); their Theorem T-3 shows that the treatment assignment models that satisfy unordered monotonicity for each pair of instrument values can be represented as an ARUM. Not all ARUM models satisfy unordered monotonicity, however; unordered monotonicity excludes a more general class of two-way flows. We will illustrate this point on one of our leading examples in Section˜3.3.
@@ 1.2 Assignment to Treatment
    where the vector $(u_{it})_{t\in\mathcal{T}}$ is distributed independently of $Z_{i}$ and is continuously distributed with full support on $\mathbb{R}^{|\mathcal{T}|}$.
@@
If we add a third treatment value so that $\mathcal{T}=\{0,1,2\}$, the ARUM assumption starts to bite as it excludes direct two-way flows in the treatment model. However, Assumption 1 is far from sufficient to identify interesting treatment effects in general. In order to better understand what is needed, we now resort to the notion of response-groups of observations, whose members share the same mapping from instruments $z$ to treatments $t$. We first state a general definition [^11] This is analogous to the definitions in Heckman and Pinto (2018)..
@@
To illustrate, consider the binary instrument/binary treatment case. It has a priori $2^{2}=4$ response vectors, $R\in\{00,01,10,11\}$ with corresponding response-groups $C_{00},C_{01},C_{10},C_{11}$. In this notation, the first number refers to a treatment value with $z=0$ and the second number with $z=1$. For instance, $C_{01}$ refers to those with $T_{i}(0)=0$ and $T_{i}(1)=1$, while the composite response-group $C_{\ast 1}$, for which $R(0)=\{0,1\}$, represents the union of $C_{01}$ and $C_{11}$ The LATE-monotonicity assumption implies that either $C_{01}$ or $C_{10}$ is empty.
@@ 2 Targeting
We start by introducing additional assumptions on the underlying treatment model. We will illustrate these assumptions on three examples: the “binary instrument model” or the $2\times T$” model; the $3\times 3$ model”; and a generalized Roy model. We first define them briefly.
@@
Suppose that agents choose occupations $t=0,\ldots,\lvert\mathcal{T}\rvert-1$ on the basis of their expected wages $w_{i}(t)=\bar{w}(t)+\eta_{it}$, net of labor disutilities $d_{i}(z,t)=\bar{d}_{z}(t)+v_{it}$ which depend on the values of the instruments:
$T_{i}(z)=\arg\max_{t=0,\ldots,\lvert\mathcal{T}\rvert-1}(w_{i}(t)-d_{i}(z,t)).$
Potential wages are $Y_{i}(t)=w_{i}(t)+\varepsilon_{it}$. We observe $Z_{i}$, the chosen occupation $T_{i}=T_{i}(Z_{i})$, and realized wages $Y_{i}=Y_{i}(T_{i})$. If the vector of variables $\{(\eta_{it},v_{it})\}_{t=0}^{\lvert\mathcal{T}\rvert-1}$ is independent of $Z_{i}$, this is an ARUM model with $U_{z}(t)=\bar{w}(t)-\bar{d}_{z}(t)$ and $u_{it}=\eta_{it}-v_{it}$.
$u_{i1}-u_{i0}$$u_{i2}-u_{i0}$$P_{z}$$T_{i}(z)=0$$T_{i}(z)=1$$T_{i}(z)=2$
@@
Under Assumption˜1, assignment to treatment is governed by the differences in mean values $(U_{z}(t)-U_{z}(\tau))$ and by the differences in unobservables $u_{it}-u_{i\tau}$. Only the former depend on the instrument. From now on, we assume that there is a reference treatment value $t_{0}$ whose mean utility does not depend on the value of the instrument:
@@
In many applications, $t=0$ is a “no-treatment” value, and instruments only change the mean utilities of the other treatments. For instance, tuition subsidies, investment credits, and invitations to training programs have no effect for those who do not attend college, do not invest, or choose not to train. Assumption˜2 seems natural in such cases [^12] In the generalized Roy model (Example 3), it holds if the disutility of occupation 0 does not depend on the values of the instruments.. For a counter-example, consider a program of unconditional cash transfers with different values $z$, for which we observe the purchases $t$ of several categories of goods the following month. If a household decides to save the transfer ($t=0$), its mean (discounted) utility will still depend on the value of the transfer $z$ that it received [^13] In that case one could define targeting with the function $\tilde{U}_{z}(t)=U_{z}(t)-U_{z}(0)$. We have not explored the consequences of this alternative definition..
@@
denote the maximum value of $U_{z}(t)$ over $z\in\mathcal{Z}$ and the set of maximizers, respectively. If $Z^{\ast}(t)$ is not all of $\mathcal{Z}$, then we will say that the instrument values $z\in Z^{\ast}(t)$ target treatment value $t$; and we write $t\in T^{\ast}(z)$. If $z$ does not target any treatment, we define $T^{\ast}(z)=\emptyset$. We denote $\mathcal{T}^{\ast}$ the set of targeted treatments and $\mathcal{Z}^{\ast}=\operatorname*{{\bigcup}}_{t\in\mathcal{T}^{\ast}}\;Z^{\ast}(t)$ the set of targeting instruments.